# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 6.1 Auxiliary arrangement I
Let $W_{\Gamma}$ be the Coxeter group of type $H_{3}$. Suppose $\Gamma$ has vertex set $\{a,b,c\}$, with $m_{ab}=5$, $m_{bc}=3$ and $m_{ac}=2$. Let $\mathcal{A}$ be the collection of reflection hyperplanes in $\mathbb{R}^{3}$ arising from the canonical representation $W_{\Gamma}\to GL(3,\mathbb{R})$.
Let $\mathsf{C}$ be the simplicial complex obtained by intersecting the unit sphere of $\mathbb{R}^{3}$ with elements in $\mathcal{A}$. In other words, $\mathsf{C}$ is the Coxeter complex associated with $W_{\Gamma}$. Each element in $\mathcal{A}$ gives a wall of $\mathsf{C}$.
@@
We define a sub-collection of walls in $\mathsf{C}$ as follows. Takes three consecutive vertices $\{\theta_{i}\}_{i=1}^{3}$ in a wall of $\mathsf{C}$ such that $\theta_{1}$ and $\theta_{3}$ are of type $\hat{c}$, and $\theta_{2}$ is type $\hat{b}$. Let $\mathcal{H}$ be the collection of walls of $\mathsf{C}$ which passes at least one of $\theta_{1},\theta_{2}$ or $\theta_{3}$. See Figure 6 left for $\mathcal{H}$. We also think $\mathcal{H}$ as a central arrangement in $\mathbb{R}^{3}$.
Let $H\subset\mathcal{H}$ be a wall passing through $\theta_{1}$. We consider the deconing of the arrangement $\mathcal{H}$ with respect to $H$. This gives arrangement $\mathcal{H}^{\prime}$ of affine hyperplanes in $\mathbb{R}^{2}$, depicted in Figure 6 right.
@@
Let $\Sigma_{\mathcal{H}^{\prime}}$ (resp. $\Sigma_{\mathcal{H}}$) be the dual polyhedron (cf. Section 2.2) associated with $\mathcal{H}^{\prime}$ (resp. $\mathcal{H}$), see Figure 7 left. Recall that there is an embedding $i_{H}:\Sigma_{\mathcal{H}^{\prime}}\to\Sigma_{\mathcal{H}}$ whose image is a maximal subcomplex of $\Sigma_{\mathcal{A}}$ which is contained in one side of $H$.
Let $\widehat{\Sigma}_{\mathcal{H}^{\prime}}$ (resp. $\widehat{\Sigma}_{\mathcal{H}}$) be the associated Salvetti complex. Let $X=\Sigma_{\mathcal{H}^{\prime}}$ and $\widehat{X}=\widehat{\Sigma}_{\mathcal{H}^{\prime}}$. Then the embedding $i_{H}:\Sigma_{\mathcal{H}^{\prime}}\to\Sigma_{\mathcal{H}}$ induces an embedding $\hat{i}_{H}:\widehat{\Sigma}_{\mathcal{H}^{\prime}}\to\widehat{\Sigma}_{\mathcal{H}}$ which is $\pi_{1}$-injective by Lemma 2.13.
@@
Let $E_{1},E_{2}$ be two edges of $X_{11}$ as in Figure 7, and let $E_{3},E_{5},E_{6}$ be the edges of $X_{13}$ in Figure 7. Let $E_{4}$ be the edge in $X_{12}$ in Figure 7. For $1\leq i\leq 6$, let $\widehat{E}_{i}$ be the associated subcomplexes of $\widehat{X}$.
@@
The group $\pi_{1}\widehat{X}$ splits as a graph of groups whose underlying graph is a linear graph with 4 vertices. The vertex groups (from left to right) are $\pi_{1}\widehat{X}_{1},\pi_{1}\widehat{X}_{2}$, $\pi_{1}\widehat{X}_{3}$ and $\pi_{1}\widehat{X}_{4}$; and the edge groups (from left to right) are $\pi_{1}\widehat{Y}_{1}$, $\pi_{1}\widehat{Y}_{2}$ and $\pi_{1}\widehat{Y}_{3}$.
@@
It suffices to show the maps $\pi_{1}\widehat{Y}_{i}\to\pi_{1}\widehat{X}_{i}$ and $\pi_{1}\widehat{Y}_{i}\to\pi_{1}\widehat{X}_{i+1}$ induced by inclusion are injective for $i=1,2,3$. We only show this for $\pi_{1}\widehat{Y}_{1}\to\pi_{1}\widehat{X}_{1}$, as the other inclusions are similar. Note that $\pi_{1}\widehat{X}_{1}$ splits as amalgamation
$\pi_{1}\widehat{X}_{11}*_{\pi_{1}A}\pi_{1}\widehat{X}_{12}$
where $A=\widehat{X}_{11}\cap\widehat{X}_{12}$, as $\pi_{1}A\to\pi_{1}\widehat{X}_{11}$ is injective (because of the retraction $\widehat{X}_{11}\to A$ as in Definition 2.9) and $\pi_{1}A\to\pi_{1}\widehat{X}_{12}$ is injective (because $\widehat{X}_{12}$ is a product of $A$ and $\widehat{X}_{12}\cap\widehat{Y}_{1}$). By [Ser02, Page 6, Proposition 3], the injectivity of $\pi_{1}\widehat{Y}_{1}\to\pi_{1}\widehat{X}_{1}$ would follow if we can show $\pi_{1}A\cap\pi_{1}(\widehat{X}_{11}\cap\widehat{Y}_{1})$ is trivial in $\pi_{1}\widehat{X}_{11}$, and $\pi_{1}A\cap\pi_{1}(\widehat{X}_{12}\cap\widehat{Y}_{1})$ is trivial in $\pi_{1}\widehat{X}_{12}$. The first statement follows by considering the retraction $\widehat{X}_{11}\to A$ which maps $\widehat{X}_{12}\cap\widehat{Y}_{1}$ to a single point, and the second statement follows from the product structure of $\widehat{X}_{12}$. ∎
@@
We now define a simplex complex of group structure $\mathcal{U}$ on $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})$ as in Figure 6.1, where the underlying complex $U$ is a union of two triangles and all the groups over 2-faces are trivial. Local groups over vertices and edges are fundamental groups of subcomplexes of $\widehat{X}$ as labeled in Figure 6.1. The morphisms between the local groups are induced by inclusions of the associated subcomplexes. By Lemma 2.13, all the morphisms between the local groups are injective, hence $\mathcal{U}$ is a simple complex of groups. Moreover, it follows from Lemma 6.2 that each local group injects into $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})$. Thus $\mathcal{U}$ is developable with $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})=\pi_{1}\mathcal{U}$.
Let $\mathbb{U}$ be the development complex of $\mathcal{U}$ (cf. [BH99, Chapter II.12]). We now give an alternative description of $\mathbb{U}$. Let $\widetilde{K}$ be the universal cover of $\widehat{X}_{1}\cup\widehat{X}_{2}$. Vertices of $\mathbb{U}$ are in 1-1 correspondence with lifts of $\widehat{X}_{ij}$ in $\widetilde{K}$ ($1\leq i\leq 2,1\leq j\leq 2$). These lifts are called standard subcomplexes of $\widetilde{K}$. Each vertex of $\mathbb{U}$ has a well-defined type, which is one of $\{\widehat{X}_{ij}\}_{1\leq i\leq 2,1\leq j\leq 2}$. Two vertices of $\mathbb{U}$ are adjacent in $\mathbb{U}$ if their associated subcomplexes of $\widetilde{K}$ have non-empty intersection, and their types correspond to adjacent vertices in $U$. Three mutually adjacent vertices of $\mathbb{U}$ form a triangle, if their associated subcomplexes have non-empty common intersection. From this description, we also see that $\mathbb{U}$ is simply connected, as $\widetilde{K}$ is simply connected.
There is a natural action $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})\curvearrowright\mathbb{U}$, with the quotient complex being $U$. We metrize $U$ such that it is made of two flat right-angled isosceles triangles, with right-angles at the vertices labeled by $\pi_{1}(\widehat{X}_{21})$ and $\pi_{1}(\widehat{X}_{12})$. This pulls back to a piecewise Euclidean metric on $\mathbb{U}$.
@@
As $\mathbb{U}$ is simply-connected, we need to the link of each vertex is CAT$(1)$. As $\widehat{X}_{12}$ splits as a product of $\widehat{X}_{11}\cap\widehat{X}_{12}$ and $\widehat{X}_{12}\cap\widehat{X}_{22}$, we know the link of a vertex of type $\widehat{X}_{22}$ is a complete bipartite graph with each edge having length $=\pi/2$. Thus this link is CAT$(1)$. Similarly the link at any vertex of type $\widehat{X}_{21}$ is CAT$(1)$.
@@
Let $\widehat{Z}=\widehat{X}_{11}\cup\widehat{X}_{22}\cup\widehat{X}_{21}$. Then there is an embedding $\widehat{Z}\to\widehat{\Sigma}_{\mathcal{A}}$ defined as follows. Consider $\widehat{Z}\subset\widehat{X}=\widehat{\Sigma}_{\mathcal{H}^{\prime}}\to\widehat{\Sigma}_{\mathcal{H}}$ where the second map is the map $\hat{i}_{H}$ after Definition 6.1. Let $c:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{H}}$ and $\hat{c}:\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{H}}$ be as in Section 2.5. Note that there is a unique face $F$ of $\Sigma_{\mathcal{A}}$ such that $c(F)=i_{H}(X_{11})$. Moreover, the collection of elements of $\mathcal{A}$ dual to $F$ is identical to the collection of elements of $\mathcal{H}$ dual to $i_{H}(X_{11})$. Thus $c$ maps $F$ homeomorphically to $i_{H}(X_{11})$, and consequently $\hat{c}$ map $\widehat{F}$ homeomorphically to $\hat{i}_{H}(\widehat{X}_{11})$. Similarly discussion applies to $\hat{i}_{H}(\widehat{X}_{22})$ and $\hat{i}_{H}(\widehat{X}_{21})$, which gives a lift $\hat{i}_{H}(\widehat{Z})$ to a subcomplex of $\widehat{\Sigma}_{\mathcal{A}}$. Let $i:\widehat{Z}\to\widehat{\Sigma}_{\mathcal{A}}$ be such embedding.
@@
Note that if $i(P)$ is null-homotopic in $\widehat{\Sigma}_{\mathcal{A}}$, then $\hat{c}\circ i(P)=\hat{i}_{H}(P)$ is null-homotopic in $\widehat{\Sigma}_{\mathcal{H}}$. Thus $P$ is null-homotopic in $\widehat{X}=\widehat{\Sigma}_{\mathcal{H}^{\prime}}$ by Lemma 2.13. Now Lemma 6.2 implies $P$ is null-homotopic in $\widehat{X}_{1}\cup\widehat{X}_{2}$. ∎
We end this section by recording the following lemma for later use. For three points $x,y,z$ in a CAT$(0)$ space, we use $\angle_{y}(x,z)$ to denote the Alexandrov between the geodesic segments $\overline{yx}$ and $\overline{yz}$ (cf. [BH99, Chapter II.3.1]).
@@ 6.2 Auxiliary arrangement II
Let $\mathsf{C}$ and $\mathcal{A}$ be as before, defined from the Coxeter group of type $H_{3}$.
@@
We define a sub-collection of walls in $\mathsf{C}$ as follows. Takes four consecutive vertices $\{\theta_{i}\}_{i=1}^{4}$ in a wall of $\mathsf{C}$ such that $\theta_{1}$ is of type $\hat{c}$, $\theta_{2},\theta_{4}$ are of type $\hat{a}$ and $\theta_{3}$ is of type $\hat{b}$. Let $\mathcal{K}$ be the collection of walls of $\mathsf{C}$ which passes at least one of $\{\theta_{i}\}_{i=1}^{4}$. See Figure 8 left for $\mathcal{K}$. We also think $\mathcal{K}$ as a central arrangement in $\mathbb{R}^{3}$.
Let $H\subset\mathcal{K}$ be the wall passing through $\theta_{1}$ as in the boundary circle of Figure 8 left6453. We consider the deconing of the arrangement $\mathcal{K}$ with respect to $H$. This gives us a planar arrangement $\mathcal{K}^{\prime}$, depicted in Figure 8 right.
@@
Let $\mathsf{Y}$ be the union of $\mathsf{X}_{31}$ together with three square faces sharing an edge with $\mathsf{X}_{31}$ (see the shaded part in Figure 9). Define $\mathsf{X}^{\prime}=\cup_{i=2}^{4}\mathsf{X}_{i}$. Let $\widehat{\mathsf{X}}_{i}$, $\widehat{\mathsf{X}}_{31}$, $\widehat{\mathsf{Y}}$, $\widehat{\mathsf{X}}^{\prime}$ be the associated subcomplex of $\widehat{\mathsf{X}}$. Note that there is an embedding $\mathsf{Y}\to\Sigma_{\mathcal{A}}$ by first embedding $\mathsf{Y}\subset\mathsf{X}$ into $\Sigma_{\mathcal{K}}$ as a maximal subcomplex in one side of $H\subset\mathcal{K}$, and lifting the image of $\mathsf{Y}$ in $\Sigma_{\mathcal{K}}$ with respect to $c:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{K}}$ in the same way as in Section 6.1. Note that the face $\mathsf{X}_{31}$ is sent to the face of $\Sigma_{\mathcal{K}}$ dual to $\theta_{2}$. This induces an embedding $i:\widehat{\mathsf{Y}}\to\widehat{\Sigma}_{\mathcal{A}}$.
@@
The map $\widehat{\mathsf{X}}^{\prime}$ to $\widehat{\mathsf{X}}$ is $\pi_{1}$-injective.
This follows from the analogue of Lemma 6.2 in the setting of $\pi_{1}\widehat{\mathsf{X}}$, which can be proved by a similar argument as Lemma 6.2.
@@
Let $\omega$ be a loop in $\widehat{\mathsf{Y}}$. If $i(\omega)$ is null-homotopic in $\widehat{\Sigma}_{\mathcal{A}}$, then $\omega$ is null-homotopic in $\widehat{\mathsf{X}}^{\prime}$.
The proof of this lemma is almost identical to Lemma 6.5, using Lemma 6.8.

### v3

@@ 6 Some sub-arrangements of the $H_{3}$-arrangement
It is natural to ask whether Falk complexes associated deconing of arrangement of type $H_{3}$ can be equipped with an equivariant non-positive curved metric, similar to what happens in Section 5.2 for type $A_{n}$. This is unfortunately not clear. However, if we are willing to consider the deconing of sub-arrangements of type $H_{3}$, then it is possible to arrange non-positive curved metric on the associated Falk subcomplex. However, it is rather subtle to decide which sub-arrangements we should use - if the sub-arrangement is too small, then it is useless for our ultimate goal, namely to understand 6-cycles in the Artin complex of type $H_{3}$; if the sub-arrangement is too large, then there might be no way to metrize the associated Falk complex to make it non-positively curved.
@@ 6.1 Auxiliary arrangement I
Let $W_{S}$ be the Coxeter group of type $H_{3}$. Suppose $S=\{a,b,c\}$, with $m_{ab}=5$, $m_{bc}=3$ and $m_{ac}=2$. Let $\mathcal{A}$ be the collection of reflection hyperplanes in $\mathbb{R}^{3}$ arising from the canonical representation $W_{S}\to GL(3,\mathbb{R})$.
Let $\mathfrak{C}$ be the simplicial complex obtained by intersecting the unit sphere of $\mathbb{R}^{3}$ with elements in $\mathcal{A}$. Then $\mathfrak{C}$ is the Coxeter complex associated with $W_{S}$. Each element in $\mathcal{A}$ gives a wall of $\mathfrak{C}$.
@@
We define a sub-collection of walls in $\mathfrak{C}$ as follows. Takes three consecutive vertices $\{\theta_{i}\}_{i=1}^{3}$ in a wall of $\mathfrak{C}$ such that $\theta_{1}$ and $\theta_{3}$ are of type $\hat{c}$, and $\theta_{2}$ is type $\hat{b}$. Let $\mathcal{H}$ be the collection of walls of $\mathfrak{C}$ which passes at least one of $\theta_{1},\theta_{2}$ or $\theta_{3}$. See Figure 5 left for $\mathcal{H}$. We also think $\mathcal{H}$ as a central arrangement in $\mathbb{R}^{3}$.
Let $H\subset\mathcal{H}$ be a wall passing through $\theta_{1}$. We consider the deconing of the arrangement $\mathcal{H}$ with respect to $H$. This gives arrangement $\mathcal{H}^{\prime}$ of affine hyperplanes in $\mathbb{R}^{2}$, in Figure 5 right.
@@
Let $\Sigma_{\mathcal{H}^{\prime}}$ (resp. $\Sigma_{\mathcal{H}}$) be the dual polyhedron (cf. Section 4.1) associated with $\mathcal{H}^{\prime}$ (resp. $\mathcal{H}$), see Figure 6 left. Recall that there is an embedding $i_{H}:\Sigma_{\mathcal{H}^{\prime}}\to\Sigma_{\mathcal{H}}$ whose image is a maximal subcomplex of $\Sigma_{\mathcal{A}}$ which is contained in one side of $H$. Let $\widehat{\Sigma}_{\mathcal{H}^{\prime}}$ (resp. $\widehat{\Sigma}_{\mathcal{H}}$) be the associated Salvetti complex. Let $X=\Sigma_{\mathcal{H}^{\prime}}$ and $\widehat{X}=\widehat{\Sigma}_{\mathcal{H}^{\prime}}$. Then the embedding $i_{H}:\Sigma_{\mathcal{H}^{\prime}}\to\Sigma_{\mathcal{H}}$ induces an embedding $\hat{i}_{H}:\widehat{\Sigma}_{\mathcal{H}^{\prime}}\to\widehat{\Sigma}_{\mathcal{H}}$ which is $\pi_{1}$-injective by Lemma 4.11.
@@
The group $\pi_{1}\widehat{X}$ splits as a graph of groups whose underlying graph is a linear graph with four vertices. The vertex groups (from left to right) are $\pi_{1}\widehat{X}_{1},\pi_{1}\widehat{X}_{2}$, $\pi_{1}\widehat{X}_{3}$ and $\pi_{1}\widehat{X}_{4}$; and the edge groups (from left to right) are $\pi_{1}\widehat{Y}_{1}$, $\pi_{1}\widehat{Y}_{2}$ and $\pi_{1}\widehat{Y}_{3}$. Moreover, the inclusion $\widehat{X}_{ij}\to\widehat{X}$ is $\pi_{1}$-injective for $1\leq i\leq 4$ and $1\leq j\leq 2$.
@@
For the first assertion of the lemma, it suffices to show the maps $\pi_{1}\widehat{Y}_{i}\to\pi_{1}\widehat{X}_{i}$ and $\pi_{1}\widehat{Y}_{i}\to\pi_{1}\widehat{X}_{i+1}$ induced by inclusion are injective for $i=1,2,3$. We only show this for $\pi_{1}\widehat{Y}_{1}\to\pi_{1}\widehat{X}_{1}$, as the other inclusions are similar. Note that $\pi_{1}\widehat{X}_{1}$ splits as amalgamation $\pi_{1}\widehat{X}_{11}*_{\pi_{1}A}\pi_{1}\widehat{X}_{12}$ where $A=\widehat{X}_{11}\cap\widehat{X}_{12}$, as $\pi_{1}A\to\pi_{1}\widehat{X}_{11}$ is injective (because of the retraction $\widehat{X}_{11}\to A$ as in Definition 4.6) and $\pi_{1}A\to\pi_{1}\widehat{X}_{12}$ is injective (because $\widehat{X}_{12}$ is a product of $A$ and $\widehat{X}_{12}\cap\widehat{Y}_{1}$). By [Ser02, Page 6, Proposition 3], the injectivity of $\pi_{1}\widehat{Y}_{1}\to\pi_{1}\widehat{X}_{1}$ would follow if we can show $\pi_{1}A\cap\pi_{1}(\widehat{X}_{11}\cap\widehat{Y}_{1})$ is trivial in $\pi_{1}\widehat{X}_{11}$, and $\pi_{1}A\cap\pi_{1}(\widehat{X}_{12}\cap\widehat{Y}_{1})$ is trivial in $\pi_{1}\widehat{X}_{12}$. The first statement follows by considering the retraction $\widehat{X}_{11}\to A$ which maps $\widehat{X}_{12}\cap\widehat{Y}_{1}$ to a single point, and the second statement follows from the product structure of $\widehat{X}_{12}$.
For the second assertion of the lemma, the previous paragraph already implies that $\widehat{X}_{11}\to\widehat{X}_{1}$ and $\widehat{X}_{1}\to\widehat{X}$ is $\pi_{1}$-injective, hence the same holds for $\widehat{X}_{11}\to\widehat{X}$. The case of $\widehat{X}_{ij}$ is similar. ∎
@@
We now define a simple complex of group structure $\mathcal{U}$ on $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})$ as in Figure 6.1, where the underlying complex $U$ is a union of two triangles and all the groups over 2-faces are trivial. Local groups over vertices and edges are fundamental groups of subcomplexes of $\widehat{X}$ as labeled in Figure 6.1. The morphisms between the local groups are induced by inclusions of the associated subcomplexes. By Lemma 4.11, all the morphisms between the local groups are injective, hence $\mathcal{U}$ is a simple complex of groups. It follows from Lemma 6.2 that each local group injects into $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})$. Thus $\mathcal{U}$ is developable with $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})=\pi_{1}\mathcal{U}$.
Let $\mathbb{U}$ be the development complex of $\mathcal{U}$ (see [BH99, Theorem II.12.18]). We now give an alternative description of $\mathbb{U}$. Let $\widetilde{K}$ be the universal cover of $\widehat{X}_{1}\cup\widehat{X}_{2}$. Vertices of $\mathbb{U}$ are in 1-1 correspondence with lifts of $\widehat{X}_{ij}$ in $\widetilde{K}$ ($1\leq i\leq 2,1\leq j\leq 2$). These lifts are called standard subcomplexes of $\widetilde{K}$. Each vertex of $\mathbb{U}$ has a well-defined type, which is one of $\{\widehat{X}_{ij}\}_{1\leq i\leq 2,1\leq j\leq 2}$. Two vertices of $\mathbb{U}$ are adjacent in $\mathbb{U}$ if their associated subcomplexes of $\widetilde{K}$ have non-empty intersection, and their types correspond to adjacent vertices in $U$. Three mutually adjacent vertices of $\mathbb{U}$ form a triangle, if their associated subcomplexes have non-empty common intersection.
There is a natural action $\pi_{1}(\widehat{X}_{1}\cup\widehat{X}_{2})\curvearrowright\mathbb{U}$, with the quotient complex being $U$. We metrize $U$ such that it is made of two flat right-angled isosceles triangles of equal size, with right-angles at the vertices labeled by $\pi_{1}(\widehat{X}_{21})$ and $\pi_{1}(\widehat{X}_{12})$. This pulls back to a piecewise Euclidean metric on $\mathbb{U}$.
@@
As $\mathbb{U}$ is simply-connected [BH99, Corollary II.12.21], we need to show the link of each vertex is CAT$(1)$. As $\widehat{X}_{12}$ splits as a product of $\widehat{X}_{11}\cap\widehat{X}_{12}$ and $\widehat{X}_{12}\cap\widehat{X}_{22}$, we know the link of a vertex of type $\widehat{X}_{22}$ is a complete bipartite graph with each edge having length $=\pi/2$. Thus this link is CAT$(1)$. Similarly the link at any vertex of type $\widehat{X}_{21}$ is CAT$(1)$.
@@
Let $\widehat{Z}=\widehat{X}_{11}\cup\widehat{X}_{22}\cup\widehat{X}_{21}$. Then there is an embedding $\widehat{Z}\to\widehat{\Sigma}_{\mathcal{A}}$ defined as follows. Consider $\widehat{Z}\subset\widehat{X}=\widehat{\Sigma}_{\mathcal{H}^{\prime}}\to\widehat{\Sigma}_{\mathcal{H}}$ where the second map is the map $\hat{i}_{H}$ after Definition 6.1. Let $\kappa:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{H}}$ and $\hat{\kappa}:\widehat{\Sigma}_{\mathcal{A}}\to\widehat{\Sigma}_{\mathcal{H}}$ be as in Section 4.4. Note that there is a unique face $F$ of $\Sigma_{\mathcal{A}}$ such that $\kappa(F)=i_{H}(X_{11})$. Moreover, the collection of elements of $\mathcal{A}$ dual to $F$ is identical to the collection of elements of $\mathcal{H}$ dual to $i_{H}(X_{11})$. Thus $\kappa$ maps $F$ homeomorphically to $i_{H}(X_{11})$, and consequently $\hat{\kappa}$ map $\widehat{F}$ homeomorphically to $\hat{i}_{H}(\widehat{X}_{11})$. Similarly discussion applies to $\hat{i}_{H}(\widehat{X}_{22})$ and $\hat{i}_{H}(\widehat{X}_{21})$, which gives a lift $\hat{i}_{H}(\widehat{Z})$ to a subcomplex of $\widehat{\Sigma}_{\mathcal{A}}$. Let $i:\widehat{Z}\to\widehat{\Sigma}_{\mathcal{A}}$ be such embedding.
@@
Note that if $i(P)$ is null-homotopic in $\widehat{\Sigma}_{\mathcal{A}}$, then $\hat{\kappa}\circ i(P)=\hat{i}_{H}(P)$ is null-homotopic in $\widehat{\Sigma}_{\mathcal{H}}$. Thus $P$ is null-homotopic in $\widehat{X}=\widehat{\Sigma}_{\mathcal{H}^{\prime}}$ by Lemma 4.11. Now Lemma 6.2 implies $P$ is null-homotopic in $\widehat{X}_{1}\cup\widehat{X}_{2}$. ∎
We end this section by recording the following lemma for later use. For three points $x,y,z$ in a CAT$(0)$ space, we use $\angle_{y}(x,z)$ to denote the Alexandrov between the geodesic segments $\overline{yx}$ and $\overline{yz}$ (cf. [BH99, Chapter II.3.1]). Let $E_{1},E_{2}$ be two edges of $X_{11}$ as in Figure 6, and let $E_{3},E_{5},E_{6}$ be the edges of $X_{13}$ in Figure 6. Let $E_{4}$ be the edge in $X_{12}$ in Figure 6. For $1\leq i\leq 6$, let $\widehat{E}_{i}$ be the associated subcomplexes of $\widehat{X}$.
@@ 6.2 Auxiliary arrangement II
Let $\mathfrak{C}$ and $\mathcal{A}$ be as before, defined from the Coxeter group of type $H_{3}$.
@@
We define a sub-collection of walls in $\mathfrak{C}$ as follows. Takes four consecutive vertices $\{\theta_{i}\}_{i=1}^{4}$ in a wall of $\mathfrak{C}$ such that $\theta_{1}$ is of type $\hat{c}$, $\theta_{2},\theta_{4}$ are of type $\hat{a}$ and $\theta_{3}$ is of type $\hat{b}$. Let $\mathcal{K}$ be the collection of walls of $\mathfrak{C}$ which passes at least one of $\{\theta_{i}\}_{i=1}^{4}$. See Figure 7 left for $\mathcal{K}$. We also think $\mathcal{K}$ as a central arrangement in $\mathbb{R}^{3}$.
Let $H\subset\mathcal{K}$ be the wall passing through $\theta_{1}$ as in the boundary circle of Figure 7 left. We consider the deconing of the arrangement $\mathcal{K}$ with respect to $H$. This gives us a planar arrangement $\mathcal{K}^{\prime}$, depicted in Figure 7 right.
@@
Let $\mathsf{Y}$ be the union of $\mathsf{X}_{31}$ together with three square faces sharing an edge with $\mathsf{X}_{31}$ (see the shaded part in Figure 8). Define $\mathsf{X}^{\prime}=\cup_{i=2}^{4}\mathsf{X}_{i}$. Let $\widehat{\mathsf{X}}_{i}$, $\widehat{\mathsf{X}}_{31}$, $\widehat{\mathsf{Y}}$, $\widehat{\mathsf{X}}^{\prime}$ be the associated subcomplex of $\widehat{\mathsf{X}}$. Note that there is an embedding $\mathsf{Y}\to\Sigma_{\mathcal{A}}$ by first embedding $\mathsf{Y}\subset\mathsf{X}$ into $\Sigma_{\mathcal{K}}$ as a maximal subcomplex in one side of $H\subset\mathcal{K}$, and lifting the image of $\mathsf{Y}$ in $\Sigma_{\mathcal{K}}$ with respect to $\kappa:\Sigma_{\mathcal{A}}\to\Sigma_{\mathcal{K}}$ in the same way as in Section 6.1. Note that the face $\mathsf{X}_{31}$ is sent to the face of $\Sigma_{\mathcal{K}}$ dual to $\theta_{2}$. This gives an embedding $i:\widehat{\mathsf{Y}}\to\widehat{\Sigma}_{\mathcal{A}}$. The next lemma can be proved in the same way as Lemma 6.2.
@@
The group $\pi_{1}\widehat{\mathsf{X}}$ splits as a graph of groups whose underlying graph is a linear graph with 4 vertices. The vertex groups (from left to right) are $\pi_{1}\widehat{\mathsf{X}}_{1},\pi_{1}\widehat{\mathsf{X}}_{2}$, $\pi_{1}\widehat{\mathsf{X}}_{3}$ and $\pi_{1}\widehat{\mathsf{X}}_{4}$. In particular, the inclusion $\widehat{\mathsf{X}}^{\prime}\to\widehat{\mathsf{X}}$ is $\pi_{1}$-injective.
The proof of the next lemma is identical to Lemma 6.5, using Lemma 6.8.
@@
Given a loop $\omega$ in $\widehat{\mathsf{Y}}$ such that $i(\omega)$ is null-homotopic in $\widehat{\Sigma}_{\mathcal{A}}$, then $\omega$ is null-homotopic in $\widehat{\mathsf{X}}^{\prime}$.